In a certain medium, E = 10 cos (108t — 3y) ax V/m. What type of medium is it?
1
Expert's answer
2020-01-10T09:45:30-0500

plane waves in lossless dielectrics

In our case, we have ω=108, β=3

we can write

"\frac{ω}{β} =\sqrt{\frac{1}{µε}}"


Answer

lossless dielectrics